# Tensor products of bimodules and bimodule pairs over monoids

### v2

@@ Abstract.
We modify the well-known tensor product of modules over a semiring, in order to treat modules over hyperrings, and, more generally, for bimodules (and bimagmas) over monoids. The tensor product of residue hypermonoids is functorial. Special attention is paid to different kinds of morphisms and the work of Nakamura and Reyes.
@@
hyperfield, hypergroup, Krasner, residue, semiring, subgroup, supertropical algebra, surpassing relation, pair, tropical.
Key words and phrases:
bimodule, hyperfield, hyperring, pair, magma, quotient hyperring, surpassing relations, tensor product, weak morphism
@@ 1. Introduction
This is part of an ongoing project in attempting to find a general algebraic framework which is suitable for tropical mathematics, and also to handle varied structures such as idempotent semirings, hyperfields, and so forth. In the process, we need to bypass negation as much as feasible. This idea can be found in [6, 7], and was implemented for blueprints in [15, 16], and put in a more general context in [24]. A minimalist set of axioms, for a “pair,” was introduced in [11], and made more precise in [2], as an associative structure $(\mathcal{A},+)$ with a distinguished substructure $\mathcal{A}_{0}$ taking the place of a zero element, which is acted on by a set $\mathcal{T}$ which often is a group. To avoid further repetition, we refer to the introductions of [1] and [11] for more historical background.
@@
Pairs can be described in terms of universal algebra, and are closed under products and sums and polynomial extensions, and provide a concrete tensor category in the sense of [9, 21]. The main example in this paper is for hypermodules. The tensor construction for residue (quotient) hypermodules is functorial.
@@
However, tensor products are not so malleable when one engages with the sorts of morphisms which arise in hyperfields, such as “weak morphisms” (Definition 2.26). [4] came up with a fully satisfactory theory for monoids, (also see [3]), but it seems that the presence of the extra operation of addition often hampers basic results such as an adjoint isomorphism theorem. One could attempt to define the tensor product of weak morphisms on simple tensors, but even this need not be well-defined without some further assumption. In this paper we do obtain weak morphisms for tensor extensions, and for tensor products when one of the tensor components is free.
@@
Recently, Nakayama and Reyes [20] have tackled the issue of tensor products of hypermagmas (in which hyperaddition $\boxplus$ is replaced by a binary operation *), in an interesting paper which also provides intriguing examples of noncommutative hypergroups. Their solution was to declare that non-simple tensors are vacuous, which is needed to obtain a general categorical tensor product of hypermagmas. The downside is that one could lose associativity of addition, as seen in Example 5.3.
@@ 1.1. Shape of the paper
We start by reviewing the basics of bimagmas, also endowed with a two-sided action, and of pairs, possibly with surpassing relations. There are three classes of morphisms (each pertaining to its own category): homomorphisms, $\preceq$-morphisms, and weak morphisms. Our main example in this paper is “hyperpairs,” inspired by hyperfield theory (and more generally the hypermagmas of [20]), although many more examples can be found in [1] and [11].
Then we bring in the main notion of this paper, the tensor product. We take a standard semiring-theoretic approach which is appropriate for the categories involving homomorphisms, but to handle the tensor product of $\preceq$-morphisms or weak morphisms in general, we need either to restrict to free bimodules, or to turn more towards ideas in [20].
As in the familiar theory of tensor products of modules, and rings, we differentiate between bimodules (resp. bimagmas) and semirings (resp. pre-semirings). The tensor extension of a pair works surprisingly well.
@@ 2.1. Underlying algebraic structures
    A magma is a set $\mathcal{A}$ with a binary operation $*:(\mathcal{A}\cup\{\infty\})\times(\mathcal{A}\cup\{\infty\})\to\mathcal{A}\cup\infty$, not necessarily associative, satisfying $\infty*b=b*\infty=\infty$ for all $b\in\mathcal{A}.$ (Thus $\infty$ is an absorbing element.) The magma is total if the operation is total, i.e., $*:\mathcal{A}\times\mathcal{A}\to\mathcal{A}$ [^1] In most applications the magma is total, but we shall utilize $\infty$ to describe tensor products. Given any total operation, one can formally adjoin the absorbing element $\infty$.
    In this paper a magma always has a neutral element, often denoted $\iota$, i.e., $\iota*b=b*\iota=b$ for all $b\in\mathcal{A}$222 In [20] these magmas are called unital. [20] calls this element $e,$ but here we use $e$ for another purpose, cf. §2.3.
@@
    A monoid is a magma with an associative operation, not necessarily commutative. An additive semigroup is an abelian monoid, with the operation denoted by “$+$,” and with a zero element $\mathbb{0}.$
@@
    A semiring $\mathcal{A}$ is a semifield if $(\mathcal{A},\cdot)$ is a group.
2.1.1. $(\mathcal{T}_{1},\mathcal{T}_{2})$-magmas and bimagmas
We take some notions from [2], slightly more generally, taking [20] into account.
@@ 2.1.1. $(\mathcal{T}_{1},\mathcal{T}_{2})$-magmas and bimagmas
    A left $\mathcal{T}$-set is a set $\mathcal{A}$ together with a (left) $\mathcal{T}$-action $\mathcal{T}\times\mathcal{A}\to\mathcal{A}$ (denoted as concatenation), for which $\mathbb{1}b=b$ and $a_{1}(a_{2}b)=(a_{1}a_{2})b$ for all $b\in\mathcal{A}$ and $a_{i}\in\mathcal{T}$.
@@
    A $(\mathcal{T}_{1},\mathcal{T}_{2})$-biset is a left $\mathcal{T}_{1}$-set and right $\mathcal{T}_{2}$-set $\mathcal{A}$, for which $(a_{1}b)a_{2}=a_{1}(ba_{2})$ for all $a_{i}\in\mathcal{T}_{i}$ and $b\in\mathcal{A}.$
@@
    A left $\mathcal{T}$-magma is a magma $(\mathcal{A},*,\iota)$, together with a (left) $\mathcal{T}$-action $\mathcal{T}\times\mathcal{A}\to\mathcal{A}$ (also denoted as concatenation), for which
    1. (a)
      $\iota$ is absorbing, i.e. $a\iota=\iota,\ \text{for all}\;a\in\mathcal{T}.$
    2. (b)
      The action is distributive over $\mathcal{T}$, in the sense that
      $a(b_{1}*b_{2})=ab_{1}*ab_{2},\quad\text{for all}\;a\in\mathcal{T},\;b_{i}\in\mathcal{A}.$
      (In particular, $a(b_{1}*b_{2})=\infty$ if and only if $ab_{1}*ab_{2}=\infty.$)
    Then we adjoin $\iota$ to $\mathcal{T}$, and denote $\mathcal{T}_{\iota}=\mathcal{T}\cup{\iota}$, declaring $\iota b=\iota$ for each $b\in\mathcal{A}$. We call the elements of $\mathcal{T}$ tangible. $\mathcal{T}$ will normally be understood from the context; in case of ambiguity, we write $\mathcal{T}_{\mathcal{A}}$ for emphasis.
@@
    A left $\mathcal{T}^{\prime}$-submagma of a left $\mathcal{T}$-magma $\mathcal{A}$ is a left submagma $\mathcal{A}^{\prime}$ over some set $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ containing $\mathbb{1}.$
@@
    A $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma $\mathcal{A}$ is a $({\mathcal{T}_{1}}_{0},{\mathcal{T}_{2}}_{0})$-biset which is both a left $\mathcal{T}_{1}$-magma and a right $\mathcal{T}_{2}$-magma.
  6. (vi)
    A $\mathcal{T}$-bimagma is a $(\mathcal{T},\mathcal{T})$-bimagma.
  7. (vii)
    $\mathcal{T}$ is $\mathcal{A}$-central if $ab=ba$ for all $a\in\mathcal{T}$ and $b\in\mathcal{A}.$
@@
    If $\mathcal{A}$ did not already contain a neutral element $\iota,$ we already could adjoin it formally by declaring its operation on all $b\in\mathcal{A}$ by $\iota*b=b*\iota=b$, and $a\iota=\iota$ for all $a\in\mathcal{T}.$
@@
    More generally, $\mathcal{T}$ could be an arbitrary set, with changes mutatis mutandis, dropping Defininition 2.2(i). For example, $\mathcal{T}$ could be, say, a set of matrix units.
(So far these definitions are rather general, and could include for example the case where $\mathcal{A}$ is a monoid with unit element $\iota$, and $\mathcal{T}$ is its set of monoid endomorphisms.)
@@
    A $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma $\mathcal{A}$ is left weakly admissible if $\mathcal{T}_{1}\subseteq\mathcal{A}$.
@@
    A left weakly admissible $(\mathcal{T}_{1},\mathcal{T}_{2})$-magma $\mathcal{A}$ is called left admissible if $\mathcal{A}$ is spanned by $\mathcal{T}_{1}$.
@@
    A right weakly admissible, resp. right admissible $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma is defined analogously to (i) and (ii).
  4. (iv)
    A weakly admissible $\mathcal{T}$-bimagma $\mathcal{A}$ is a left and right weakly admissible $(\mathcal{T},\mathcal{T})$-bimagma (identifying both copies of $\mathcal{T}$ inside $\mathcal{A}$).
@@
Let $\mathcal{A}$ be a weakly admissible $\mathcal{T}$-bimagma. Then $\langle\mathcal{T}\rangle$ (the submagma generated by $\mathcal{T}$) is an admissible $\mathcal{T}$-submagma of $\mathcal{A},$ which is $\mathcal{A}$ if and only if $\mathcal{A}$ is left admissible.
@@ 2.2. Prepairs and pairs
A $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair $(\mathcal{A},\mathcal{A}_{0})$ (sometimes written $(\mathcal{A},\mathcal{A}_{0};*)$ if we want to specify the operation *) is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma $(\mathcal{A},*,\iota)$ together with a $(\mathcal{T}_{1},\mathcal{T}_{2})$-sub-bimagma $\mathcal{A}_{0}$ i.e., $a_{1}b_{0}\in\mathcal{A}_{0}$ and $b_{0}a_{2}\in\mathcal{A}_{0}$ for all $a_{i}\in\mathcal{T}_{i}$ and $b_{0}\in\mathcal{A}_{0}.$
@@
    A $\mathcal{T}$-pair $(\mathcal{A},\mathcal{A}_{0})$ is a weakly admissible $\mathcal{T}$-prepair, where $\mathcal{A}_{0}$ is closed under the operation *, i.e., $\mathcal{A}_{0}$ is a left $\mathcal{T}_{1}$-submagma of $\mathcal{A}$ [^3] In [1] $\mathcal{A}_{0}$ is not required to be closed under ()(*)( ∗ ). In previous work [1, 2, 11] we assumed that all pairs are proper, to dismiss the degenerate case of $\mathcal{T}=\mathcal{A}_{0}=\mathcal{A}.$ But this is precisely the case treated so successfully in [10], so we permit it here. We suppress $\mathcal{T}$ in the notation when it is understood.
@@
    A $\mathcal{T}$-pair $(\mathcal{A},\mathcal{A}_{0})$ is said to be proper if $\mathcal{A}_{0}\cap\mathcal{T}=\emptyset$.
  3. (iii)
    A $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair $(\mathcal{A},\mathcal{A}_{0})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma prepair when $\mathcal{A}$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma.
  4. (iv)
    A gp-pair is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-pair for which $\mathcal{T}_{1},\mathcal{T}_{2}$ are groups.
2.3. Property N [1, §3.1]
For $b\in\mathcal{A}$, an element $b^{\dagger}$ is a pseudo-negative of $b$ if $b+b^{\dagger}=b^{\dagger}+b\in\mathcal{A}_{0}$, and $b^{\dagger}\in\mathcal{T}_{\iota}$ when $b\in\mathcal{T}_{\iota}$. Note that a pseudo-negative need not be unique.
@@
We say that a pair $(\mathcal{A},\mathcal{A}_{0})$ satisfies weak Property N if $\mathbb{1}$ has a pseudo-negative $\mathbb{1}^{\dagger}$, with $\mathbb{1}^{\dagger}\mathcal{T}\subseteq\mathcal{T}$ and $\mathbb{1}^{\dagger}b=b\mathbb{1}^{\dagger}$ for each $b\in\mathcal{A}$. In this case, fixing $\mathbb{1}^{\dagger}$, we define $e:=\mathbb{1}*\mathbb{1}^{\dagger}\in\mathcal{A}_{0}.$ Then, we denote $b^{\dagger}=b\mathbb{1}^{\dagger}$, and $b^{\circ}=b*b^{\dagger}=b^{\dagger}*b$, for all $b\in\mathcal{A}.$ Let $\mathcal{A}^{\circ}=\{b^{\circ}:b\in\mathcal{A}\},$ and $\mathcal{T}^{\circ}=\{a^{\circ}:a\in\mathcal{T}\}.$
@@
    $\mathcal{T}^{\circ}=\{a^{\circ}:a\in\mathcal{T}\}$.
@@
  4. (iv)
    $(\mathcal{A},\mathcal{A}_{0})$ is $\circ$-distributive if $(a_{1}*a_{2})e=a_{1}e*a_{2}e$ for all $a_{i}\in\mathcal{T}.$
  5. (v)
    $(\mathcal{A},\mathcal{A}_{0})$ is $\circ$-idempotent if $e*e=e.$
By [2, Lemma 3.9], if $\mathcal{A}$ is a $\mathcal{T}$-magma, then the action of (iii) makes $\mathcal{A}$ a $\mathcal{T}\cup\mathcal{T}^{\circ}$-magma, when $(\mathcal{A},\mathcal{A}_{0})$ is $\circ$-distributive and $\circ$-idempotent.
Example 2.10.
  1. (i)
    The pair $(R,\{\mathbb{0}\})$, where $R$ is as in Remark 2.3(i), is called the classical pair.
  2. (ii)
    The classical field pair is the classical gp-pair $(F,\{\mathbb{0}\})$, where $F$ is a field.
Important Note 2.11.
Philosophically, $\mathcal{A}_{0}$ takes the place of $\mathbb{0}$ (or, multiplicatively, $\mathbb{1}$) in classical mathematics. The significance is that since modules need not have negation (for example, $\mathbb{N}$), $\mathbb{0}$ has no significant role except as a place marker in linear algebra.

### v4

@@ Abstract.
We modify the well-known tensor product of modules over a semiring, in order to treat modules over hyperrings, and, more generally, for bimodules (and semi-algebras) over monoids. The tensor product of residue hypermodules is functorial with respect to this construction. Special attention is paid to different kinds of morphisms and the work of Nakamura and Reyes.
@@
bimodule, hyperfield, hyperring, pair, magma, quotient hyperring, residue hypermodule, surpassing relations, tensor product, weak morphism
@@ 1. Introduction
This is part of an ongoing project to find a general algebraic framework that is suitable for tropical mathematics, and also to handle varied structures such as idempotent semirings, F1-geometry/algebra, hyperfields, and so forth. In the process, we bypass negation as much as feasible. This idea originally is found in [6, 7] (also see [3]), and was implemented for blueprints in [15, 16], and put in a more general context in [24]. A minimalist set of axioms, for a “pair,” was introduced in [11], and made more precise in [2], as an additive semigroup $(\mathcal{A},+)$ together a distinguished null subsemigroup $\mathcal{A}_{0}$, taking the place of a zero element, both of which are acted on by a set $\mathcal{T}$ which often is a group. To avoid further repetition, we refer to the introductions of [1] and [11] for more background. The main example in this paper comes from the class of hypermodules.
@@
Pairs can be studied by means of universal algebra, and are closed under products, sums, and polynomial extensions, and provide a concrete tensor category with respect to homomorphisms, in the sense of [9, 21]. An early reference for tensor products is [13].
However, tensor products are not so malleable when one engages with the sorts of morphisms that arise in hyperfields, such as “weak morphisms” (Definition 2.24) and $\preceq$-morphisms” (Definition 2.18). To begin with, it is not even clear what we want to take the tensor product over: the hyperfield or its power set? Since the power set of a hyperfield need not be a semiring, the latter seems an impossible task.
Remark 1.1.
Let us start with the observation that the classical treatment of the tensor product of a right module and a left module over a ring $C$ (e.g., see [23, Chapter 18]) only uses the fact that $C$ is a monoid. So we shall talk of tensor products over a multiplicative monoid.
@@
A fully satisfactory tensor theory for tensor products of monoids over a monoid is provided in [4], but it seems that the presence of the extra operation of addition often gets in the way. One could attempt to define the tensor product of weak morphisms via simple tensors, but even this need not be well-defined without further assumptions.
@@
Recently, Nakayama and Reyes [20] have tackled the issue of tensor products in full generality for hypermagmas (in which the hyperaddition $\boxplus$ is replaced by a binary operation *), in an interesting paper which also provides intriguing examples of noncommutative hypergroups. Their solution was to declare that non-simple tensors are vacuous, which is needed to obtain a general categorical tensor product of hypermagmas. The downside is that the classical theory of tensor products of algebras is not a special case of their results. Furthermore, one may lose associativity of addition, as seen in Example 5.3, and basic constructions such as polynomials cannot be viewed in terms of tensor products.
Our goal in this paper is to see how far one can carry a theory that includes the familiar tensor product of modules, which is compatible with the residue (quotient) hypermodule construction originated by Krasner [14].
When using homomorphisms for our morphisms, we get the analog of the familiar module-theoretic results in Theorem 4.24 and its corollaries; also see Remark 5.17 for the adjoint correspondence.
But weak morphisms are more relevant for hypermodules We do obtain weak morphisms of tensor extensions, and for tensor products when one of the tensor components is free, such as for polynomials, and get a chunk of the adjoint isomorphism theorem in Lemma 5.20.
In order to be applicable to hyperfields, the theory must not require distributivity of multiplication over addition, since the power sets of certain hyperfields are not distributive, and then taking polynomials ruins associativity of multiplication. Thus of necessity, as in [20], for a general theory one is led to magmas, dropping associativity in multiplication, and [20] frames the theory over magmas. Nevertheless, after a brief excursion to magmas, we focus on modules over monoids.
@@ 1.1. Shape of the paper
We start by reviewing the basics of magmas, also endowed with a two-sided action by a set, and of “pairs,” possibly with a surpassing relation denoted as $\preceq$. There are three major classes of morphisms (each pertaining to its own category): homomorphisms (the morphisms in universal algebra), $\preceq$-morphisms (which preserve the surpassing relation), and weak morphisms (which preserve the connection to the null subsemigroup). Our main motivation in this paper is from “hyperpairs,” inspired by hyperfield theory (and more generally the hypermagmas of [20]), which include supertropical algebra, although many more examples of pairs can be found in [2] and [11].
Then we bring in the main notion of this paper, the tensor product. We start with a standard module-theoretic approach which is appropriate for categories involving homomorphisms. The definition of tensor product in Definition 4.2 is standard. Some care needs to be taken for defining tensor products of homomorphisms, in Theorems 4.12 and 4.24].
However, more care is needed when dealing with $\preceq$-morphisms or weak morphisms:
  • We can restrict to free modules, as done in Theorem 4.24 when one of the morphisms is a homomorphism; this permits us to extend $\preceq$-morphisms or weak morphisms on tensor extensions.
  • We may modify ideas espoused in [20], in Proposition 5.11.
  • There is a natural tensor product map for hypermodules (Theorem 5.13), but which is not resonant with category theory.
We obtain tensor products of $\preceq$-morphisms in Theorem 5.13.
As in the familiar theory of tensor products of modules and algebras, we also obtain the $\mathcal{T}$-semialgebra version of tensor extensions in Theorem 5.15.
Finally we turn to a pillar of the category theory of modules, the adjoint isomorphism theorem. This seems impossible to obtain in full, but a partial result is given in Theorem 5.18. The closest we get to an adjoint isomorphism theorem involving weak morphisms is Theorem 5.19.
@@ 2.1. Underlying algebraic structures
    A magma is a set $\mathcal{A}$ with a binary operation $(\mathcal{A}\cup\{\infty\})\times(\mathcal{A}\cup\{\infty\})\to\mathcal{A}\cup\{\infty\}$, not necessarily associative, satisfying $\infty*b=b*\infty=\infty$ for all $b\in\mathcal{A}.$ (Thus $\infty$ is an absorbing element.) The magma is total if the operation is total, i.e., $\mathcal{A}\times\mathcal{A}\to\mathcal{A}$. [^1] In most applications the magma is total, but we shall utilize $\infty$ to describe tensor products. Given any total operation, one can formally adjoin the absorbing element $\infty$.
    In this paper a magma always has a neutral element, often denoted $\iota$, i.e., $\iota*b=b*\iota=b$ for all $b\in\mathcal{A}$. [^2] In [20] these magmas are called unital, with the neutral element designated as $e,$ but we have used $e$ for another purpose, cf. §2.2.1.
@@
    A monoid is a magma with an associative operation, not necessarily commutative. An additive semigroup is a commutative monoid, with the operation denoted by “++,” endowed with a zero element $\mathbb{0}.$
@@
    A semiring $\mathcal{A}$ is a semifield if $(\mathcal{A}\setminus\{\mathbb{0}\},\cdot)$ is a group.
2.1.1. $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagmas
We recall some notions from [2], slightly more generally, taking [20] into account.
@@ 2.1.1. $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagmas
    A left $\mathcal{T}$-magma is a magma $(\mathcal{A},*,\iota)$, together with a (left) $\mathcal{T}$-action $\mathcal{T}\times\mathcal{A}\to\mathcal{A}$ (also denoted as concatenation), for which, for all $a,a_{i}\in\mathcal{T}$, $b,b_{i}\in\mathcal{A}$,
    1. (a)
      $\mathbb{1}b=b$.
    2. (b)
      $a_{1}(a_{2}b)=(a_{1}a_{2})b$.
    3. (c)
      $\iota$ is absorbing, i.e. $a\iota=\iota.$
    4. (d)
      The action is distributive over $\mathcal{T}$, in the sense that $a(b_{1}*b_{2})=ab_{1}*ab_{2}.$ (In particular, $a(b_{1}*b_{2})=\infty$ if and only if $ab_{1}=\infty$ or $ab_{2}=\infty.$)
    Then we adjoin an absorbing element $\iota_{\mathcal{T}}$ to $\mathcal{T}$, and denote $\mathcal{T}_{\iota}=\mathcal{T}\cup\{\iota_{\mathcal{T}}\}$, declaring $\iota_{\mathcal{T}}b=\iota$ for each $b\in\mathcal{A}$. We call the elements of $\mathcal{T}$ tangible. $\mathcal{T}$ will normally be understood from the context; in case of ambiguity, we write $\mathcal{T}_{\mathcal{A}}$ for emphasis.
@@
    A left $\mathcal{T}$-submagma of a left $\mathcal{T}$-magma $\mathcal{A}$ is a left submagma closed under the $\mathcal{T}$-action.
@@
    A $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma $\mathcal{A}$ is a left $\mathcal{T}_{1}$-magma and a right $\mathcal{T}_{2}$-magma satisfying $(a_{1}b)a_{2}=a_{1}(ba_{2})$ for all $a_{i}\in\mathcal{T}_{i}$ and $b\in\mathcal{A}.$
@@
    A $\mathcal{T}$-bimagma is a $(\mathcal{T},\mathcal{T})$-bimagma.
@@
    $\mathcal{T}$ is central in a $\mathcal{T}$-bimagma $\mathcal{A}$ if $ab=ba$ for all $a\in\mathcal{T}$ and $b\in\mathcal{A}.$
@@
    More generally, $\mathcal{T}$ could be a partial monoid, i.e., with multiplication only defined on certain pairs of elements, with changes mutatis mutandis, using Defininition 2.2(i)(b) only when the right side is defined. For example, $\mathcal{T}$ could be, say, a set of matrix units with multiplication defined only on compatible matrix units.
@@
    If a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma $\mathcal{A}$ did not already contain a neutral element $\iota,$ we could adjoin it formally by declaring its operation on all $b\in\mathcal{A}$ by $\iota*b=b*\iota=b$, and $a_{i}\iota=\iota a_{i}=\iota$ for all $a\in\mathcal{T}.$
(So far these definitions are rather general, and could include for example the case where $\mathcal{A}$ is a monoid with unit element $\iota$, and $\mathcal{T}$ is its semigroup of monoid endomorphisms.)
@@
    A left $\mathcal{T}$-magma $\mathcal{A}$ is weakly admissible if $\mathcal{T}\subseteq\mathcal{A}$.
@@
    A weakly admissible, resp. admissible, right $\mathcal{T}$-bimagma is defined analogously to (i) and (ii).
@@
    A weakly admissible $\mathcal{T}$-bimagma $\mathcal{A}$ is a left and right weakly admissible $\mathcal{T}$-magma (identifying both copies of $\mathcal{T}$ inside $\mathcal{A}$).
@@
Let $\mathcal{A}$ be a weakly admissible $\mathcal{T}$-magma. Then $\langle\mathcal{T}\rangle$ (the submagma generated by $\mathcal{T}$) is an admissible $\mathcal{T}$-submagma of $\mathcal{A},$ which is $\mathcal{A}$ if and only if $\mathcal{A}$ is admissible.
@@ 2.2. Prepairs and pairs
A $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair $(\mathcal{A},\mathcal{A}_{0})$ (sometimes written $(\mathcal{A},\mathcal{A}_{0};*)$ if we want to specify the operation *) is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma $(\mathcal{A},*,\iota)$ together with a $(\mathcal{T}_{1},\mathcal{T}_{2})$-sub-bimagma $\mathcal{A}_{0}$ (i.e., $a_{1}b_{0}\in\mathcal{A}_{0}$ and $b_{0}a_{2}\in\mathcal{A}_{0}$ for all $a_{i}\in\mathcal{T}_{i}$ and $b_{0}\in\mathcal{A}_{0}).$
A $(\mathcal{T},\mathcal{T})$-prepair is also called a $\mathcal{T}$-prepair. We suppress $\mathcal{T}$ in the notation when it is understood.
@@
    A pair $(\mathcal{A},\mathcal{A}_{0})$ is a weakly admissible prepair with $ab=ba$ for each $a\in\mathcal{T},$ $b\in\mathcal{A}$. [^3] In [1] $\mathcal{A}_{0}$ was not required to be closed under ()(*). In previous work [1, 2, 11] we assumed that all pairs are proper, to dismiss the degenerate case of $\mathcal{T}=\mathcal{A}_{0}=\mathcal{A}.$ But this is precisely the case treated so successfully in [10], so we permit it here.
@@
    A pair $(\mathcal{A},\mathcal{A}_{0})$ is said to be proper if $\mathcal{A}_{0}\cap\mathcal{T}=\emptyset$.
Important Note 2.8.
Philosophically, $\mathcal{A}_{0}$ takes the place of $\mathbb{0}$ in classical mathematics. The reason is that since modules need not have negation (for example, $\mathbb{N}$), $\mathbb{0}$ has no significant role except as a place marker in linear algebra.
2.2.1. Property N [1, §3.1]
@@
We say that a pair $(\mathcal{A},\mathcal{A}_{0})$ satisfies weak Property N if there is an element $\mathbb{1}^{\dagger}\in\mathcal{T}$ with $\mathbb{1}*\mathbb{1}^{\dagger}=\mathbb{1}^{\dagger}*\mathbb{1}\in\mathcal{A}_{0}$. (But $\mathbb{1}^{\dagger}$ need not be unique.) In this case, fixing $\mathbb{1}^{\dagger}$, we define $e:=\mathbb{1}*\mathbb{1}^{\dagger}\in\mathcal{A}_{0}.$ Then we denote $b^{\dagger}=b\mathbb{1}^{\dagger}$, and $b^{\circ}=b*b^{\dagger}=b^{\dagger}*b$, for all $b\in\mathcal{A}.$
Note for $a\in\mathcal{T}$ that $a^{\circ}=a(\mathbb{1}*\mathbb{1}^{\dagger})=ae\in\mathcal{A}_{0}.$
@@
    $\mathcal{T}^{\circ}:=\{a^{\circ}:a\in\mathcal{T}\}\subseteq\mathcal{A}_{0}$.
@@
    $(\mathcal{A},\mathcal{A}_{0})$ is $\circ$-distributive if $(b_{1}*b_{2})e=b_{1}e*b_{2}e$ for all $b_{i}\in\mathcal{A}.$
  4. (iv)
    $(\mathcal{A},\mathcal{A}_{0})$ is $\mathcal{A}_{0}$-idempotent if bb=bb*b=b for all $b\in\mathcal{A}_{0}.$
  5. (v)
@@
By [2, Lemma 3.9], the action of (v) makes $\mathcal{A}$ a $\mathcal{T}\cup\mathcal{T}^{\circ}$-bimagma.